# Cut-and-paste for impulsive gravitational waves with $\Lambda$: The mathematical analysis

### v1

@@ 3.4 Associated geodesics
Now to formulate a precise result we first establish a notation for the limiting geodesics. Clearly in front of the impulse, that is for $\lambda<0$ corresponding to $U_{\varepsilon}<0$, $\gamma_{\varepsilon}$ will converge to the seed geodesic $\gamma^{-}$. Similarly, behind the impulse, that is for $\lambda>0$ corresponding to $U_{\varepsilon}>0$, $\gamma_{\varepsilon}$ will also converge to a geodesic $\gamma^{+}$ of the background (A)dS space. Here $\gamma^{+}=(U^{+},V^{+},Z_{p}^{+})$ is specified by the values of $\gamma_{\varepsilon}$ and $\dot{\gamma}_{\varepsilon}$ upon leaving the regularization zone: Indeed it is shown in the course of the proof of Theorem 3.1 (cf. [43, eq. (34) and below]) that there is a parameter value $\beta_{\varepsilon}>0$ such that $U_{\varepsilon}(\beta_{\varepsilon})=\varepsilon$ and that $\beta_{\varepsilon}\searrow 0$ for $\varepsilon\to 0$, c.f. [45, Lem. A2]. Moreover, the corresponding values of $\gamma_{\varepsilon}(\beta_{\varepsilon})$ and $\dot{\gamma}_{\varepsilon}(\beta_{\varepsilon})$ converge (cf. [45, Prop. 5.3]). More precisley, we have
@@
Note that the limiting geodesics $\tilde{\gamma}$ of (3.17) can be interpreted as the geodesics of the distributional spacetime (2.6), (2.7). Keep in mind, however, that (3.17) does not solve the (formal) geodesic equations of the distributional spacetime (see [45, Eq. (2.6)], [36, Eq. (28)]) by the lack of a consistent solution concept. Indeed the low regularity of $\tilde{\gamma}$ does not allow to insert it into these equations.
@@ 4.1 The null geodesic generators & the ‘discontinuous transformation’
This family of null geodesics is parameterized by three real constants fixing the positions at the parameter value ${\mathcal{U}}=\lambda=0$, i.e., $\gamma^{g}_{4D}(0)=(0,{\mathcal{V}}_{0},x_{0},y_{0}$).
Next we write the null generators (4.1) in the five-dimensional representation of Section 2.3 but still parameterized by the 4D-data $({\mathcal{V}}_{0},x_{0},y_{0})$ (cf. [39, Eq. (26)])
@@
$\!\!\!\!\gamma_{4D}[{\mathcal{V}}_{0},x_{0},y_{0}]({\mathcal{U}})=\left(\begin{array}[]{c}{\mathcal{U}}\\ {\mathcal{V}}_{0}+\Theta({\mathcal{U}})\,{\mathcal{H}}^{\mathrm{i}}+{\mathcal{U}}_{+}\,\frac{1}{2}\big{(}({\mathcal{H}}_{,x}^{\mathrm{i}})^{2}+({\mathcal{H}}_{,y}^{\mathrm{i}})^{2}\big{)}\\ x_{0}^{j}+{\mathcal{U}}_{+}{\mathcal{H}}_{,j}^{\mathrm{i}}\\ \end{array}\right)\!,$ (4.9)
@@
We have hence formally recovered the ‘discontinuous transformation’ from a special family of global limiting null geodesics, which can be interpreted as the geodesics of the distributional spacetime (2.6), (2.7), cf. the one before last paragraph of Section 3.4.
Moreover, the behavior of these geodesics can be vividly depicted, see Figure 2 (cf. [39, Fig. 2], and [40, Fig. 6]) in a way that directly generalizes Penrose’s original illustration for the $\Lambda=0$-case in [31, Fig. 2]: The null geodesic generators of (A)dS starting in the ‘lower half’ $(A)d\mathcal{S}^{-}$ (i.e., for $U=\lambda<0$) due to their interaction with the wave impulse do not continue as unbroken null generators into $(A)d\mathcal{S}^{+}$ (indicated by the dashed line in the upper left parts). Rather they jump at $\{U=\lambda=0\}$, cf. the $\Theta$-term in (4.9) (hence in (4.13)) but also get refracted, cf. the $\lambda_{+}$-terms, to become the appropriate null generators of $(A)d\mathcal{S}^{+}$.
Figure 2: The null geodesic generators of (A)dS and their interaction with the impulisve wave.
The upshot is that these ‘broken geodesic generators’ basically are the coordinate lines of the coordinate system in which the metric becomes continuous, i.e., (2.2). But we do not only have these limiting geodesics at our hand but also the generalized geodesics of Theorem 3.1. This will allow us to geometrically regularize the transformation, which we will explicitly do next.
4.2 The geometrically regularized transformation
Using the ideas laid out above we now give the explicit form of the transformation in nonlinear generalized functions. To begin with, note that we only have carried out the nonlinear distributional analysis of the geodesics in the $5D$-form [^6] As explained in [38, p. 3], the geodesic equations in the $4D$-form are even wilder and a direct approach seems to be out of reach.. Therefore we split up the transformation in the following way: starting from the ‘continuous’ $4D$-coordinates $(u,v,Z)\equiv(u,v,X,Y)$ of (2.2) we first use the transformation (2.9) [^7] Observe that (2.9) literally transforms the conformally flat coordinates $({\mathcal{U},\mathcal{V},\eta})$ to the $5D$-coordinates. But here we transform data ‘in front’ of the wave, where $({\mathcal{U},\mathcal{V},\eta})$ are trivially related to $(u,v,Z)$, cf. (2.4) and so in (2.9) we have to replace $({\mathcal{U},\mathcal{V},\eta})$ with $(u,v,Z)$. to go to the $5D$-coordinates $(U,V,Z_{p})$. Then we use the regularised geodesics to transform
@@ 4.2 The geometrically regularized transformation
where [^8] We here write the regularisation parameter as a superscript rather than as a subscript to make the notation more appealing. $[(\gamma^{\varepsilon}_{5D}[V,Z_{p}])_{\varepsilon}]$ are the generalized solutions of the geodesic equations provided by Theorem 3.1 with data constructed from the seed geodesic $\gamma^{g}_{5D}$ of (4.3), i.e., the null generator with data $(V,Z_{p},\dot{V},\dot{Z}_{p})$ as in (4.4), but now derived from $(v,X,Y)$ instead of $({\mathcal{V}}_{0},x_{0},y_{0})$. We will specify this data explicitly below but first we turn to the final final part of the transformation. For this we use the inverse of (2.9), i.e.,
@@ 5 Analysis of the regularised transformation
In this section we finally establish that the geometrically regularized transformation $(T_{\varepsilon})_{\varepsilon}$ is a representative of a generalized diffemorphism $T=[(T_{\varepsilon})_{\varepsilon}]$ in the sense of nonlinear distributional geometry and thus give a precise mathematical meaning to the physical equivalence of the distributional and the continuous form of the metric.
@@
Definition 5.1 (Generalized diffeomorphism).
Let $\Omega\subseteq\mathbb{R}^{n}$ be open. We call $T\in{\mathcal{G}}[\Omega,\mathbb{R}^{n}]$ a generalized diffeomorphism if there exists $\eta>0$ such that
@@
We will show that $T$ is a generalized diffeomorphism and we will split up this task in two subsections.
5.1 $T$ as a generalized function
To begin with, we have to establish that the regularised transformation $T_{\varepsilon}$ gives rise to a c-bounded generalized function on $\mathbb{R}^{5}$, more precisely that $[(T_{\varepsilon})_{\varepsilon}]\in{\mathcal{G}}[\mathbb{R}^{5},\mathbb{R}^{5}]$. Recall that by (4.32) we have $T_{\varepsilon}(U,V,Z_{p})=\gamma^{\varepsilon}_{5D}[V,Z_{p}](U)$ and that so far we have only considered $\gamma^{\varepsilon}_{5D}$ as a function of $U$. Indeed, Theorem 3.1 guarantees that $[(\gamma^{\varepsilon}_{5D})_{\varepsilon}]\in{\mathcal{G}}[\mathbb{R},\mathbb{R}^{5}]$, but now we have to additionally deal with the dependence of $\gamma^{\varepsilon}_{5D}[V,Z_{p}](U)$ on $V$ and $Z_{p}$.
@@ 5.1 $T$ as a generalized function
Denoting by normal-$\nabla$ any of the derivatives $\partial_{V}$ and $\partial_{Z_{p}}$ we have in the regularisation strip $-\varepsilon\leq r\leq\beta_{\varepsilon}$ and $(V,Z_{p})$ varying in a compact set
@@
Denoting by normal-$\nabla$ any of the derivatives $\partial_{V}$ and $\partial_{Z_{p}}$ we have in the regularisation strip $-\varepsilon\leq r\leq\beta_{\varepsilon}$ and $(V,Z_{p})$ varying in a compact set that for any $n$ there is $N$ such that
@@
We proceed by induction. Clearly Lemma 5.5 provides the basis of induction. So assume that we have $\nabla^{n}(\bar{U}_{\varepsilon},\dot{\bar{U}}_{\varepsilon},\bar{Z}_{p\varepsilon},\dot{\bar{Z}}_{p\varepsilon})=O(\varepsilon^{-M})$ for some $M$. We aim for again a nested Gronwall argument for the highest order derivatives. Staring with $\bar{U}_{\varepsilon}$, and $\dot{\bar{U}}_{\varepsilon}$, we find using the integral representation (4.40)
@@
simply collecting all lower order terms in the final $O(\varepsilon^{-N})$-estimate. Observe, that the (critical) term involving $\delta_{\varepsilon}(\bar{U}_{\varepsilon})$ does not produce any (high) inverse powers of $\varepsilon$ in the highest order terms $\nabla^{n+1}(\bar{U}_{\varepsilon},\dot{\bar{U}}_{\varepsilon},\bar{Z}_{p\varepsilon})$.
@@
The transformation $[(T_{\varepsilon})_{\varepsilon}]$ is moderate and hence an element of ${\mathcal{G}}[\mathbb{R}^{5},\mathbb{R}^{5}]$.
@@ 5.2 $T$ as generalized diffeomorphism
$(U,V,Z_{p})\mapsto[T_{\varepsilon}](U,V,Z_{p})=[\gamma^{\varepsilon}_{5D}][V,Z_{p}](U)\,\in{\mathcal{G}}[R^{5},\mathbb{R}^{5}]\,,$ (5.34)
with its components given explicitly by (4.40) - (4.43) gives rise to a locally invertible generalized function $T=[(T_{\varepsilon})_{\varepsilon}]$ on some open set containing the impulsive surface. Hence we will call it a generalised diffeomorphism or generalised coordinate transformation. To do so we will extend the results of the $\Lambda=0$-case of [20] and, in particular, its more mathematically structured presentation in [10]. In fact, inspired by [10] we will decompose the transformation in a convenient way by splitting $T_{\varepsilon}$ into a ‘singular’ and a ‘convergent’ part.
@@
Outside the regularization strip $\{\alpha_{\varepsilon}\leq U\leq\beta_{\varepsilon}\}$ the transformation defaults to a smooth coordinate transform independent of $\varepsilon$ and hence possesses all the properties needed in the following arguments. Thus we may restrict ourselves to the regularization zone and there $\bar{U}_{\varepsilon}=O(\varepsilon)$ holds. As for $T_{\varepsilon}$, being able to decompose its $V$ component into $\tilde{V}_{\varepsilon}+h_{\varepsilon}$ lets us more easily compute its Jacobian. First, we focus on the regularization strip $\{\alpha_{\varepsilon}\leq U\leq\beta_{\varepsilon}\}$:
@@
Thus, all the principal minors are of the form $1+O(\varepsilon)$, and thus, in particular, $|\det(DT_{\varepsilon})|\geq\varepsilon^{N}$ for all $N\in\mathbb{N}$. Consequently, $(T_{\varepsilon}^{-1})_{\varepsilon}$ is moderate, and from $T_{\varepsilon}^{-1}\circ T_{\varepsilon}=\mathrm{id}$ we conclude that $T_{\varepsilon}^{-1}$ is c-bounded (on the image of $T_{\varepsilon}$).
In conclusion, this gives that $T=[(T_{\varepsilon})_{\varepsilon})]$ is a generalized diffeomorphism. We are, however, interested in the overall transformation (4.30), i.e., the precomposition of $T$ with (2.9) and the postcomposition with (4.15). Since they both are (classical) smooth diffeomorphsims on their respective domains we only need to observe that such a composition clearly is a generalized diffeomorphism. So in total we have:
@@
The discontinuous coordinate transform (4.30) is a generalized diffeomorphism.

### v2

@@ 3.4 Associated geodesics
Now to formulate a precise result we first establish a notation for the limiting geodesics. Clearly in front of the impulse, that is for $\lambda<0$ corresponding to $U_{\varepsilon}<0$, $\gamma_{\varepsilon}$ will converge to the seed geodesic $\gamma^{-}$. Similarly, behind the impulse, that is for $\lambda>0$ corresponding to $U_{\varepsilon}>0$, $\gamma_{\varepsilon}$ will also converge to a geodesic $\gamma^{+}$ of the background (A)dS space. Here $\gamma^{+}=(U^{+},V^{+},Z_{p}^{+})$ is specified by the values of $\gamma_{\varepsilon}$ and $\dot{\gamma}_{\varepsilon}$ upon leaving the regularisation zone: Indeed it is shown in the course of the proof of Theorem 3.1 (cf. [43, eq. (34) and below]) that there is a parameter value $\beta_{\varepsilon}>0$ such that $U_{\varepsilon}(\beta_{\varepsilon})=\varepsilon$ and that $\beta_{\varepsilon}\searrow 0$ for $\varepsilon\to 0$, c.f. [45, Lem. A2]. Moreover, the corresponding values of $\gamma_{\varepsilon}(\beta_{\varepsilon})$ and $\dot{\gamma}_{\varepsilon}(\beta_{\varepsilon})$ converge (cf. [45, Prop. 5.3]). More precisley, we have
@@
Note that the limiting geodesics $\tilde{\gamma}$ of (3.17) can be interpreted as the geodesics of the distributional spacetime (2.6), (2.7). Keep in mind, however, that (3.17) does not solve the (formal) geodesic equations of the distributional spacetime (see [45, Eq. (2.6)], [36, Eq. (28)]) by the lack of a consistent solution concept. Indeed the low regularity of $\tilde{\gamma}$ does not allow one to insert it into these equations.
@@ 4.1 The null geodesic generators & the ‘discontinuous transformation’
This family of null geodesics is parameterised by three real constants fixing the positions at the parameter value ${\mathcal{U}}=\lambda=0$, i.e., $\gamma^{g}_{4D}(0)=(0,{\mathcal{V}}_{0},x_{0},y_{0}$).
Next we write the null generators (4.1) in the five-dimensional representation of Section 2.3 but still parameterised by the 4D-data $({\mathcal{V}}_{0},x_{0},y_{0})$ (cf. [39, Eq. (26)])
@@
$\!\!\!\!\gamma_{4D}[{\mathcal{V}}_{0},x_{0},y_{0}]({\mathcal{U}})=\left(\begin{array}[]{c}{\mathcal{U}\vspace{2.0mm}}\\ {\mathcal{V}}_{0}+\Theta({\mathcal{U}})\,{\mathcal{H}}^{\mathrm{i}}+{\mathcal{U}}_{+}\,\frac{1}{2}\big{(}({\mathcal{H}}_{,x}^{\mathrm{i}})^{2}+({\mathcal{H}}_{,y}^{\mathrm{i}})^{2}\big{)}\vspace{2.0mm}\\ x_{0}^{j}+{\mathcal{U}}_{+}{\mathcal{H}}_{,j}^{\mathrm{i}}\\ \end{array}\right)\!,$ (4.9)
@@
We have hence formally recovered the ‘discontinuous transformation’ from a special family of global limiting null geodesics, which can be interpreted as the geodesics of the distributional spacetime (2.6), (2.7), cf. the penultimate paragraph of Section 3.4.
Moreover, the behaviour of these geodesics can be vividly depicted, see Figure 2 (cf. [39, Fig. 2], and [40, Fig. 6]) in a way that directly generalizes Penrose’s original illustration for the $\Lambda=0$-case in [31, Fig. 2]: The null geodesic generators of (A)dS starting in the ‘lower half’ $(A)d\mathcal{S}^{-}$ (i.e., for $U=\lambda<0$) due to their interaction with the wave impulse do not continue as unbroken null generators into $(A)d\mathcal{S}^{+}$ (indicated by the dashed line in the upper left parts). Rather they jump at $\{U=\lambda=0\}$, cf. the $\Theta$-term in (4.9) (hence in (4.13)) but also get refracted, cf. the $\lambda_{+}$-terms, to become the appropriate null generators of $(A)d\mathcal{S}^{+}$.
Figure 2: The null geodesic generators of (A)dS and their interaction with the impulsive wave.
The upshot is that these ‘broken geodesic generators’ basically are the coordinate lines of the coordinate system in which the metric becomes continuous, i.e., (2.2). But we do not only have these limiting geodesics at hand but also the generalised geodesics of Theorem 3.1. This will allow us to geometrically regularise the transformation, which we will explicitly do next.
4.2 The geometrically regularised transformation
Using the ideas laid out above we now give the explicit form of the transformation in nonlinear generalised functions. To begin with, note that we only have carried out the nonlinear distributional analysis of the geodesics in the $5D$-form [^6] As explained in [38, p. 3], the geodesic equations in the $4D$-form are even wilder and a direct approach seems to be out of reach.. Therefore we split up the transformation in the following way: starting from the ‘continuous’ $4D$-coordinates $(u,v,Z)\equiv(u,v,X,Y)$ of (2.2) we first use the transformation (2.9) [^7] Observe that (2.9) literally transforms the conformally flat coordinates $({\mathcal{U},\mathcal{V},\eta})$ to the $5D$-coordinates. But here we transform data ‘in front’ of the wave, where $({\mathcal{U},\mathcal{V},\eta})$ are trivially related to $(u,v,Z)$, cf. (2.4) and so in (2.9) we have to replace $({\mathcal{U},\mathcal{V},\eta})$ with $(u,v,Z)$. to go to the $5D$-coordinates $(U,V,Z_{p})$. Then we use the regularised geodesics to transform
@@ 4.2 The geometrically regularised transformation
where [^8] We here write the regularisation parameter as a superscript rather than as a subscript to make the notation more appealing. $[(\gamma^{\varepsilon}_{5D}[V,Z_{p}])_{\varepsilon}]$ are the generalised solutions of the geodesic equations provided by Theorem 3.1 with data constructed from the seed geodesic $\gamma^{g}_{5D}$ of (4.3), i.e., the null generator with data $(V,Z_{p},\dot{V},\dot{Z}_{p})$ as in (4.4), but now derived from $(v,X,Y)$ instead of $({\mathcal{V}}_{0},x_{0},y_{0})$. We will specify this data explicitly below but first we turn to the final final part of the transformation. For this we use the inverse of (2.9), i.e.,
@@ 5 Analysis of the regularised transformation
In this section we finally establish that the geometrically regularised transformation $(T_{\varepsilon})_{\varepsilon}$ is a representative of a generalised diffemorphism $T=[(T_{\varepsilon})_{\varepsilon}]$ in the sense of nonlinear distributional geometry and thus give a precise mathematical meaning to the physical equivalence of the distributional and the continuous form of the metric.
@@
Definition 5.1 (generalised diffeomorphism).
Let $\Omega\subseteq\mathbb{R}^{n}$ be open. We call $T\in{\mathcal{G}}[\Omega,\mathbb{R}^{n}]$ a generalised diffeomorphism if there exists $\eta>0$ such that
@@
We will show that $T$ is a generalised diffeomorphism and we will split up this task in two subsections.
5.1 $T$ as a generalised function
To begin with, we have to establish that the regularised transformation $T_{\varepsilon}$ gives rise to a c-bounded generalised function on $\mathbb{R}^{5}$, more precisely that $[(T_{\varepsilon})_{\varepsilon}]\in{\mathcal{G}}[\mathbb{R}^{5},\mathbb{R}^{5}]$. Recall that by (4.32) we have $T_{\varepsilon}(U,V,Z_{p})=\gamma^{\varepsilon}_{5D}[V,Z_{p}](U)$ and that so far we have only considered $\gamma^{\varepsilon}_{5D}$ as a function of $U$. Indeed, Theorem 3.1 guarantees that $[(\gamma^{\varepsilon}_{5D})_{\varepsilon}]\in{\mathcal{G}}[\mathbb{R},\mathbb{R}^{5}]$, but now we have to additionally deal with the dependence of $\gamma^{\varepsilon}_{5D}[V,Z_{p}](U)$ on $V$ and $Z_{p}$.
@@ 5.1 $T$ as a generalised function
Denoting by $\nabla$ any of the derivatives $\partial_{V}$ and $\partial_{Z_{p}}$ we have in the regularisation strip $-\varepsilon\leq r\leq\beta_{\varepsilon}$ and $(V,Z_{p})$ varying in a compact set
@@
We have in the regularisation strip $-\varepsilon\leq r\leq\beta_{\varepsilon}$ and for $(V,Z_{p})$ varying in a compact set
@@
Denoting by $\nabla$ any of the derivatives $\partial_{V}$ and $\partial_{Z_{p}}$ we have in the regularisation strip $-\varepsilon\leq r\leq\beta_{\varepsilon}$ and $(V,Z_{p})$ varying in a compact set that for any $n$ there is $N$ such that
@@
We proceed by induction. Clearly Lemma 5.5 provides the basis of induction. So assume that we have $\nabla^{n}(\bar{U}_{\varepsilon},\dot{\bar{U}}_{\varepsilon},\bar{Z}_{p\varepsilon},\dot{\bar{Z}}_{p\varepsilon})=O(\varepsilon^{-M})$ for some $M$. We again aim for a nested Gronwall argument for the highest order derivatives. Staring with $\bar{U}_{\varepsilon}$, and $\dot{\bar{U}}_{\varepsilon}$, we find using the integral representation (4.40)
@@
simply collecting all lower order terms in the final $O(\varepsilon^{-N})$-estimate. Observe that the (critical) term involving $\delta_{\varepsilon}(\bar{U}_{\varepsilon})$ does not produce any (high) inverse powers of $\varepsilon$ in the highest order terms $\nabla^{n+1}(\bar{U}_{\varepsilon},\dot{\bar{U}}_{\varepsilon},\bar{Z}_{p\varepsilon})$.
@@
The net of transformations $(T_{\varepsilon})_{\varepsilon}$ is moderate and hence $[(T_{\varepsilon})_{\varepsilon}]$ is an element of ${\mathcal{G}}[\mathbb{R}^{5},\mathbb{R}^{5}]$.
@@ 5.2 $T$ as generalised diffeomorphism
$(U,V,Z_{p})\mapsto[T_{\varepsilon}](U,V,Z_{p})=[\gamma^{\varepsilon}_{5D}][V,Z_{p}](U)\,\in{\mathcal{G}}[\mathbb{R}^{5},\mathbb{R}^{5}]\,,$ (5.34)
with its components given explicitly by (4.40) - (4.43) gives rise to a locally invertible generalised function $T=[(T_{\varepsilon})_{\varepsilon}]$ on some open set containing the impulsive surface. Hence we will call it a generalised diffeomorphism or generalised coordinate transformation. To do so we will extend the results of the $\Lambda=0$-case of [20] and, in particular, its more mathematically structured presentation in [10]. In fact, inspired by [10] we will decompose the transformation in a convenient way by splitting $T_{\varepsilon}$ into a ‘singular’ and a ‘convergent’ part.
@@
Outside the regularisation strip $\{\alpha_{\varepsilon}\leq U\leq\beta_{\varepsilon}\}$ the transformation defaults to a smooth coordinate transform independent of $\varepsilon$ and hence possesses all the properties needed in the following arguments. Thus we may restrict ourselves to the regularisation zone and there $\bar{U}_{\varepsilon}=O(\varepsilon)$ holds. As for $T_{\varepsilon}$, being able to decompose its $V$ component into $\tilde{V}_{\varepsilon}+h_{\varepsilon}$ lets us more easily compute its Jacobian. First, we focus on the regularisation strip $\{\alpha_{\varepsilon}\leq U\leq\beta_{\varepsilon}\}$:
@@
Thus, all the principal minors are of the form $1+O(\varepsilon)$, and hence, in particular, $|\det(DT_{\varepsilon})|\geq\varepsilon^{N}$ for some $N\in\mathbb{N}$. Consequently, $(T_{\varepsilon}^{-1})_{\varepsilon}$ is moderate, and from $T_{\varepsilon}^{-1}\circ T_{\varepsilon}=\mathrm{id}$ we conclude that $T_{\varepsilon}^{-1}$ is c-bounded (on the image of $T_{\varepsilon}$).
In conclusion, this gives that $T=[(T_{\varepsilon})_{\varepsilon})]$ is a generalised diffeomorphism. We are, however, interested in the overall transformation (4.30), i.e., the precomposition of $T$ with (2.9) and the postcomposition with (4.15). Since they both are (classical) smooth diffeomorphsims on their respective domains we only need to observe that such a composition clearly is a generalised diffeomorphism. So in total we have:
@@
The discontinuous coordinate transform (4.30) is a generalised diffeomorphism.